Find the general solution.

11. r(t) = (1 -3)<< (t)
Answer #1

Given :- 2 x(t)= •5 X(t) 1 - 2 let, D x (t): 1x (t) Ixact xi(t) X(t) = (rice xz(t) X, (t) X, () * (t) X2 (t) 2-X -5 -> eigeFor X=-i; + i -5 - -2 ti -27 reduced row echelon form is now ) solving -2+1 ( (. ])-(3) (=] => If ₂=t -t (2-3) I ti solution

Kindly comment for queries, if any and upvote

if you like it.